OpenStudy (anonymous):

if 2 +iroot3 is a root of x^2 + px + q =0 whre p, q are real then (p,q)=?

OpenStudy (shubhamsrg):

i'll give a tip,,for any polynomial like this of any degree,,it a+ib is its solution,, then 110% a-ib will also be its solution.. try now..

OpenStudy (amistre64):

i was gonna say conjugate, but that works too :)

OpenStudy (shubhamsrg):

:D

OpenStudy (anonymous):

did nt understand lol

OpenStudy (anonymous):

see if 2+iroot3 is a root, then the other root is 2-iroot3 so sum of roots = p/1=2+iroot3+2-iroot3 p=4

OpenStudy (anonymous):

got it thxxx for ur help

OpenStudy (apoorvk):

Complex roots always occur as conjugate surds. That is, if one of the roots is a+ib, then the other root will be a-ib So, you can find out the second root using this funda, which would be 2 - i*sqrt3

OpenStudy (anonymous):

and q= (2+1root3)(2-iroot3) try now, nd p will be -4 not 4

OpenStudy (apoorvk):

Now sum of roots = 4 = -p and product of roots = 2^2 + 3^2 = q = 13